Only 5 out of 1000 people get sick from a specific disease. A test for the disease is 99% accurate. This test also has a 2% false positive result (2% of the people that test positive are actually not sick with the disease). What is the probability that if you receive a positive test result, that you actually have the disease?
Answer #1

Solution:- The probability that if you receive a positive test result, that you actually have the disease is 0.1992

P(Disease) = 5/1000 = 0.005

P(Not Disease) = 1- 0.005 = 0.995